Determine the critical value for a 95% confidence interval when the sample size is 11 for the t‑distribution. Enter the positive critical value rounded to 3 decimal places.
Answer #1

Solution :

Given that,

sample size = n = 11

Degrees of freedom = df = n - 1 = 11 - 1 = 10

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,10 = 2.228

critical value = 2.228